Jose has 19 new magazines to read. Let M be the number of magazines he would have left to read after reading of them. Write an equation M relating R to . Then use this equation to find the number of magazines he would have left to read after reading 11 of them


1
Expert's answer
2022-03-27T16:04:34-0400

M = Number of unread magazines

R= Magazines already read

we need to find number of unread magazines (M) after reading 11 magazines (i.e. R = 11)


Unread magazines + magazines already read = Total number of new magazines

M + R = 19

Subtracting R from both sides

M + R - R = 19- R

M = 19 - R ..............required equation.

substituting R= 11

M = 19 - 11

M = 8

So there will be 8 magazines left to read.